OpenStudy (anonymous):

Algebra 1 question: What is the sum

OpenStudy (anonymous):



OpenStudy (michele_laino):

hint: the least common multiple between 2b and 2, is 2b. is it ok?

OpenStudy (anonymous):

what do you mean by "is it ok"

OpenStudy (abdullah1995):

lol

OpenStudy (michele_laino):

if you also think that the least common multiple is 2b

OpenStudy (anonymous):

oh, if i had to guess i wouldve said 2

OpenStudy (michele_laino):

do you think that the least common multiple is 2b?

OpenStudy (anonymous):

but i also really suck at algebra so im not sure

OpenStudy (anonymous):

maybe im thinking least common factor...

OpenStudy (anonymous):

yeah i would say 2b for LCM

OpenStudy (michele_laino):

I think taht the least common multiuple is 2b since we have to pick all common factors and uncommon factors with the highest exponent. Now our factors are: 2 and b so we have to pick both, then we have: least common multiple = 2*b=2b

OpenStudy (michele_laino):

do you remember that rule?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

now, in order to sum your fraction, we can write this: \[\Large \frac{1}{{2b}} + \frac{b}{2} = \frac{{1 + b \times b}}{{2b}}\]

OpenStudy (michele_laino):

fractions*

OpenStudy (anonymous):